OpenStudy (anonymous):

the density of ice is 910kg/m^3, that of water is 1000kg/m^3. Which part of the volume V of a block of ice is underwater? a)0.91V b)0.09V c)0.01V d)0.1V

OpenStudy (anonymous):

force of gravity=buyoncy force so, force of gravity=910*v*g and byoncy force=1000*V*g gives V=.091v

OpenStudy (anonymous):

@keshavsarda, i don't get you working: force of gravity and buoyancy=mg

OpenStudy (anonymous):

@ajprincess and @Tushara

OpenStudy (anonymous):

sorry cant help u with this, i think keshavsarda is right

OpenStudy (anonymous):

he is but I don't understand him..

OpenStudy (experimentx):